A rectangular yard measuring 37ft by 62ft is bordered (and surrounded) by a fence. Inside, a walk that is 4ft wide goes all the way along the fence. Find the area of this walk. Be sure to include the correct unit in your answer.








Answer by ankor@dixie-net.com(22740)

A rectangular yard measuring 37ft by 62ft is bordered (and surrounded) by a fence.
Inside, a walk that is 4ft wide goes all the way along the fence.
Find the area of this walk.
:
Find the overall area: 37 * 62 = 2294 sq/ft
:
The Area inside the walk dimensions, will be the overall dimensions reduced by 8 ft when there is 4 ft walkway
:
29 * 54 = 1566 sq/ft, area inside the walkway
:
Overall area - inside area = walkway area
2294 - 1566 = 728 sq/ft is the area of the walkway